OpenStudy (anonymous):

9 + (negative10) + (negative20) = A. 39 B. negative39 C. 21 D. negative21

OpenStudy (anonymous):

9+(-10)+(-20)=9-10-20=9-30=-21 D.negative 21

OpenStudy (zzr0ck3r):

write -10 for negative10

OpenStudy (anonymous):

thanks guys!!

OpenStudy (anonymous):

np